# PlaceFormer: Transformer-based Visual Place Recognition using Multi-Scale Patch Selection and Fusion

### v1

@@ V-B Latency and Memory
For real-world scalability, the VPR systems need to be efficient in terms of the time taken to extract the features and match those features. Table V-B shows the computational time and memory requirements for the state-of-the-art methods and PlaceFormer for a single query image. SP-SuperGlue, DELG, Patch-NetVLAD, and TransVPR all adopt RANSAC for their matching process, a strategy similar to PlaceFormer. However, PlaceFormer exhibits superior efficiency, being $18.4$, $12.3$, $148.4$, and $5$ times faster in terms of extraction latency, and $7.11$, $32.76$, $6.95$, and $2.9$ times faster in terms of matching latency compared to these methods. The enhanced speed of PlaceFormer in matching can be attributed to its approach of selecting key patches using attention scores, resulting in a more focused set of points for which the homography needs to be computed during RANSAC. This optimization leads to a significant reduction in computation time. $R^{2}$ormer requires a similar extraction time as that of PlaceFormer, but it needs less time for matching due to the use of transformer blocks for the matching process.
TABLE II: Time taken for extracting features, matching descriptors, and memory requirements for the descriptor for state-of-the-art methods and PlaceFormer. The latency is measured on an NVIDIA RTX 3090 GPU. For global retrieval methods, matching time and memory requirements are negligible and hence denoted with a ‘–’.
Method
Extraction
Latency (ms)
Matching
Latency (s)
Memory
NetVLAD [2]
17
SFRS [29]
203
SP-SuperGlue [30]
166
7.83 1.93
DELG [13]
197
36.04 0.37
Patch-NetVLAD [6]
1336
7.65 44.14
TransVPR [21]
45
3.19 1.17
$R^{2}$ormer [22]
9
0.3 0.5
PlaceFormer (Ours)
9
1.1 1.07

### v2

@@ V-B Latency and Memory
Table II shows the computational time and memory requirements for the state-of-the-art methods and PlaceFormer for a single query image. SP-SuperGlue, DELG, Patch-NetVLAD, and TransVPR all adopt RANSAC for their matching process, a strategy similar to PlaceFormer. However, PlaceFormer exhibits superior efficiency, being $18.4$, $12.3$, $148.4$, and $5$ times faster in terms of extraction latency, and $7.11$, $32.76$, $6.95$, and $2.9$ times faster in terms of matching latency compared to these methods. The enhanced speed of PlaceFormer in matching can be attributed to its approach of selecting key patches using attention scores, resulting in a more focused set of points for which the homography needs to be computed during RANSAC. This optimization leads to a significant reduction in computation time. MixVPR though requires the least extraction time, the method extracts a vector of size $4096$ which makes the global retrieval a tedious task. $R^{2}$ormer requires a similar extraction time as that of PlaceFormer, but it needs less time for matching due to the use of transformer blocks for the matching process.
TABLE II: Time taken for extracting features, matching descriptors, and memory requirements for the descriptor for state-of-the-art methods and PlaceFormer. For global retrieval methods, matching time and memory requirements are negligible and hence denoted with a ‘–’.
{tblr}
row1 = c, cell22 = c, cell23 = c, cell24 = c, cell32 = c, cell33 = c, cell34 = c, cell42 = c, cell43 = c, cell44 = c, cell52 = c, cell53 = c, cell54 = c, cell62 = c, cell63 = c, cell64 = c, cell72 = c, cell73 = c, cell74 = c, cell82 = c, cell83 = c, cell84 = c, cell92 = c, cell93 = c, cell94 = c, cell102 = c, cell103 = c, cell104 = c, hline1-2,10-11 = -,
Method & Extraction
Latency (ms) Matching
Latency (s) Memory
NetVLAD [2] 17 – –
SFRS [27] 203 – –
MixVPR [29] 6 – –
SP-SuperGlue [30] 166 7.83 1.93
DELG [12] 197 36.04 0.37
Patch-NetVLAD [5] 1336 7.65 44.14
TransVPR [19] 45 3.19 1.17
$R^{2}$ormer [20] 9 0.3 0.5
PlaceFormer (Ours) 9 1.1 1.07
@@
V-C Ablation Study
We perform multiple ablation experiments to further affirm the design choices made in PlaceFormer.
Patch Sizes. To assess the efficacy of employing fused patches and explore various combinations, we conduct ablations on different patch sizes and their amalgamations. In Table III, we present the performance of PlaceFormer using patches of varying sizes and combinations for re-ranking. Here, $P_{K}$ represents key patches selected from the transformer’s patch tokens. $P_{2K}$ and $P_{3K}$ denote fused key patches obtained through average pooling with kernel sizes of $2$ and $3$. When using patches of different sizes independently for re-ranking, non-fused patches $P_{K}$ yield the best results. This outcome can be attributed to the abundance of key patches and their correspondence to smaller regions, facilitating precise homography estimation.
The utilization of patches at multiple scales during correspondence estimation resulted in improved performance, contingent on the combination of patches employed. Specifically, combining non-fused key patches $P_{K}$ with the first level of fused patches $P_{2K}$, denoted as $P_{K}$ $P_{2K}$ in Table III, yielded increased recall values compared to using patches of the same size independently. This suggests that the synergistic use of key patches at different scales contributes positively to the overall performance, highlighting the effectiveness of incorporating multi-scale information for correspondence estimation in PlaceFormer.
Figure 3: Visualization of global retrieval and re-ranking results (red box- incorrect retrieval and orange box - correct retrieval); the attention maps and key patches for query and closest retrieved image at multiple scales; and the inliers estimated across patches of various scales.
TABLE III: Ablation study on various patch sizes and their combination on MSLS validation.
{tblr}
width = colspec = Q[702]Q[73]Q[73]Q[88], cell12 = c, cell13 = c, cell14 = c, cell22 = c, cell23 = c, cell24 = c, cell32 = c, cell33 = c, cell34 = c, cell42 = c, cell43 = c, cell44 = c, cell52 = c, cell53 = c, cell54 = c, cell62 = c, cell63 = c, cell64 = c, cell72 = c, cell73 = c, cell74 = c, cell92 = c, cell93 = c, cell94 = c, cell102 = c, cell103 = c, cell104 = c, cell112 = c, cell113 = c, cell114 = c, hline1-2,12 = -, Patch Size & R@1 R@5 R@10
$P_{K}$ 87.3 93.4 94.6
$P_{2K}$ 86.9 92.7 93.1
$P_{3K}$ 84.3 91.8 92.4
$P_{K}$ $P_{2K}$ 88.4 93.9 94.9
$P_{K}$ $P_{3K}$ 88.1 93.5 94.7
$P_{K}$ $P_{2K}$ $P_{3K}$ 85.4 92.7 93.1
$P_{K}$ + $P_{K}$ $P_{2K}$ 88.9 93.8 95.1
$P_{K}$ + $P_{K}$ $P_{2K}$ + $P_{2K}$ $P_{3K}$ 89.9 94.3 95.4
$P_{K}$ + $P_{K}$ $P_{2K}$ + $P_{K}$ $P_{2K}$ $P_{3K}$ 87.9 93.5 94.9
$P_{K}$ + $P_{K}$ $P_{3K}$ + $P_{K}$ $P_{2K}$ $P_{3K}$ 86.4 91.4 93.3
Further, we aggregated the number of inliers estimated through different combinations of patch sizes to assess which combination yields the most effective results for re-ranking. Through experiments, we found that utilizing the sum of inliers estimated using non-fused key patches $P_{K}$; non-fused key patches $P_{K}$ with the first level of fused patches $P_{2K}$ ($P_{K}$ $P_{2K}$); and non-fused key patches $P_{K}$ with the second level of fused patches $P_{3K}$ ($P_{K}$ $P_{3K}$), denoted as $P_{K}$ + $P_{K}$ $P_{2K}$ + $P_{K}$ $P_{3K}$ in Table III gave the best recall values. This combination of key patches is set as default and used for all the experiments.
Furthermore, the simultaneous utilization of patches of all three sizes ($P_{K}$ $P_{2K}$ $P_{3K}$) generally resulted in a decrease in recall values. This suggests that combining patches of more than two sizes for correspondence estimation may not be well-suited for optimal performance. Additionally, patches with a size exceeding that of $P_{3K}$ are not considered, as such patches may correspond to a significantly large area in the image and may not be optimal for homography computation.
TABLE IV: Ablation study on the threshold for patch selection on MSLS validation.
{tblr}
width = colspec = Q[310]Q[173]Q[173]Q[210], columneven = c, column3 = c, hline1-2,11 = -, Threshold, $\tau$ R@1 R@5 R@10
0.008 85.3 91.2 92.1
0.007 86.9 92.7 93.1
0.009 88.7 93.2 95.1
0.01 89.9 94.3 95.4
0.011 89.1 93.9 94.8
0.012 88.2 93.1 93.8
0.015 86.4 91.4 92.5
0.02 86.2 91.2 92.0
0.05 84.2 90.9 91.9
Key Patch Selection Threshold. We eliminate patch tokens and fused patches associated with potentially non-informative image regions, excluding them from the correspondence estimation. Employing a threshold parameter, $\tau$, only patches with an attention score exceeding $\tau$ contribute to the correspondence estimation process. In Table IV, we present ablation results, assessing the impact of different $\tau$ values on MSLS validation. Our experiments reveal an increase in recall values with increasing $\tau$, peaking at $\tau=0.01$. However, further increases in $\tau$ diminish performance, as non-informative patches are inadvertently included in the correspondence estimation. Ultimately, we select $\tau=0.01$ as the default value for all experiments based on its optimal balance between increased recall and avoiding the inclusion of non-informative patches.
Different Backbones. DINOv2 [31] stands out as a prominent Vision Foundational Model (VFM), proficient in tackling various vision challenges in its pre-initialized state. In Table V, we explored different variants of DINOv2 as the backbone for extracting global features and patch tokens, while using our method for re-ranking. Comparative results were obtained with DINOv2 backbones which underscore the scalability of our re-ranking approach across other backbones. Notably, the performance of DINOv2 backbones was primarily influenced by attention scores that were not fine-tuned for place recognition tasks.
TABLE V: Ablation study on MSLS Validation with various backbone architectures.
{tblr}
width = colspec = Q[465]Q[44]Q[123]Q[123]Q[146], columnodd = fg=black, column3 = c, column4 = c,fg=black, column5 = c, cell42 = c, cell52 = c, hline1-2,6 = -, Backbone Architecture & R@1 R@5 R@10
DINOv2 ViT-S/14 80.2 83.4 85.9
DINOv2 ViT-B/14 85.4 89.2 92.7
DINOv2 ViT-L/14 87.7 90.6 93.0
PlaceFormer (Ours) 89.9 94.3 95.4
V-D Visualization
In Fig. 3, we present a detailed case illustrating various retrieved images along with the attention maps and the key patches used for the retrieval process. Fig. 3 (a) shows a query image along with the top-3 retrieval using global retrieval (Fig. 3 (b)) and the top-3 results following re-ranking (Fig. 3 (c)). It can be seen that all of the top-3 retrieval using global descriptors are incorrect, while the re-ranking using multi-scale patches yields correct top-3 retrieval. Figs. 3 (d) and (f) provide insight into the attention mechanism, showcasing the attention map $A_{L}$ from the final layer of the transformer, along with the fused attention maps $A_{2L}$ and $A_{3L}$ for the query image and the closest matching reference image. The attention maps reveal that higher scores are assigned to task-relevant regions in the image like buildings, while dynamic objects like cars, cyclists, and the sky receive lower attention values. Figs 3 (e) and (g) illustrate the identification of key patches at various scales based on attention scores the query image and the closest matching reference image. The inliers estimated between these patches of different scales are depicted in Figs 3 (h), (i), and (j). The inliers illustrate that employing patches of multiple scales facilitates the identification of more correspondences compared to using patches of similar scales. The additional inliers, in turn, contribute to the improved re-ranking of images.
VI Conclusion
This paper introduces PlaceFormer, a novel approach to place recognition employing the vision transformer. PlaceFormer leverages patch tokens extracted from the vision transformer, synthesizing patches of multiple scales through fusion. The amalgamation of these multi-scale patches yields superior image retrieval results compared to existing state-of-the-art methods, which typically operate with patches of similar sizes across diverse benchmark datasets. Notably, the incorporation of attention scores from the vision transformer enables the identification of task-relevant regions in the image. Consequently, only patches corresponding to these pertinent regions are retained, effectively reducing memory usage in PlaceFormer. The selective use of key patches further accelerates the correspondence estimation for re-ranking, contributing to the overall efficiency of the proposed approach. Nevertheless, our approach has certain limitations, particularly in matching latency when contrasted with methods employing neural networks for re-ranking. The computational intensity arises from estimating homography using RANSAC. As future research, we intend to explore the development of network models capable of efficiently matching patches of various scales in diverse combinations, hence enhancing the overall efficiency of the method.